OpenStudy (anonymous):

A function is defined by these values... For which constants (a and b) is this function continuous? Where f(x) = x^3 for x <-1 ax+b for -1 _1

OpenStudy (anonymous):

So I would assume that \[\lim_{x \rightarrow 1-}\] would be \[x^{3}\]

OpenStudy (anonymous):

and \[\lim_{x \rightarrow 1+}\] would be \[x^{2}+2\]

OpenStudy (anonymous):

and \[\lim_{x \rightarrow 0-}\] would be the same as \[\lim_{x \rightarrow 0+}\] which would be ax+b. But where would I go next?

OpenStudy (freckles):

\[(-1)^3=a(-1)+b \\ a(1)+b=(1)^2+2\] you have a system of equations to solve

OpenStudy (freckles):

I set left limit as x approaches -1 equal to right limit as x approaches -1 I set left limit as x approaches 1 equal to right limit as x approaches 1

OpenStudy (freckles):

you can simplify those equations so they don't look so junky

OpenStudy (anonymous):

Okay thank you so much, this is my second time taking this course and I have this condition where I keep getting sick, so I just wanted to say that I really appreciate the extra help. People like you keep me positive and working toward my goal

OpenStudy (freckles):

glad I could help